OpenStudy (selena2345):

A Goblin Miner earns 1 cent the first day, 2 cents the second day 4 cents the third day, 8 cents the fourth day and so on doubling the amount each day. How much will the miner earn in 7 days ? 2 Weeks? 1 Month ( Make An Estimate Before Solving) Show/ Explain All Work

OpenStudy (anonymous):

7 days is easy 1+2+4+8+16+32+64

OpenStudy (anonymous):

you might notice that this is 127, one less than 2 times 64

OpenStudy (anonymous):

you have \(1, 2, 2^2, 2^3, 2^4, ...\)

OpenStudy (selena2345):

Are your answering the question right

OpenStudy (anonymous):

in 7 days you will have earned \(2^7-1=\$1.27\)

OpenStudy (selena2345):

Are your answering the question right

OpenStudy (anonymous):

in two weeks you will have \(2^{4}-1\) pennies, a calculator gives \(\$163.83\)

OpenStudy (anonymous):

I want that job if I get to work the whole month!

OpenStudy (anonymous):

me too in one month of 31 days you will have made \(2^{31}-1\) pennies, more money than i will earn in several life times

OpenStudy (selena2345):

can u help cliffSedge

OpenStudy (selena2345):

yes cliffsedge thank you

OpenStudy (anonymous):

If you have a calculator, you can try entering '1' then hitting x2 a bunch of times (counting how many times you do so.)

OpenStudy (anonymous):

Of course, that will only get you the per day earnings. You'll have to add up all the previous days' to get the total.

OpenStudy (anonymous):

But satellite's formula is a lot quicker. ;-)

OpenStudy (selena2345):

thank you agent47

OpenStudy (agent47):

Let x be the amt u earn per day... Then, in 1 day: 1cent = 2^0 cents 2 days: 2 cents = 2^1 cents 3 days: 4 cents = 2^2 cents 4 days: 8 cents = 2^3 cents... So, in n days, you get 2^(n-1) cents

OpenStudy (agent47):

I think, satellite's made a mistake.. but that's IMPOSSIBRU!

OpenStudy (agent47):

Thus, in 7 days: 2^(7-1)=2^6=64 cents

OpenStudy (agent47):

woah waitttt

OpenStudy (agent47):

That is the amount he earned on the 1st, 2nd, 3rd, 4th, etc days, and on the 7th day he earns 64 cents

OpenStudy (selena2345):

is the answer to the question finish

OpenStudy (agent47):

Not yet... the formula is: